The question says: Solve the Logarithmic equation:
2 log x - 2 = log(x-25)

Answer by stanbon(48508)

The question says: Solve the Logarithmic equation:
2 log x - 2 = log(x-25)
---------------------
log(x^2) - log(x-25) = 2
---
log[x^2/(x-25)] = 2
---
[x^2/(x-25)] = 100
---
x^2 = 100x - 2500
---
x^2 - 100x + 2500 = 0
(x-50)^2 = 0
x = 50
====================
Cheers,
Stan H.